How to Recruit Members for the League

Richard Stallman

(February 5, 1991)

The League for Programming Freedom has got off to a good start; We have around 470 members as of today. But in order to impress officials that we represent substantial public opinion, we need thousands of members.

Since August I have developed an easy and effective way of recruiting more members. I have probably found over a hundred members with this method, most of them either bare acquaintances or complete strangers. But if I do this alone, the League will at best grow linearly, and that is not enough. We need to grow exponentially, and that depends on your help.

The obvious way to look for more members is to post annoucements, but this does not work very well. We have tried many of them, and our experience shows that announcements, either electronic or physical, are not nearly as effective as one-on-one conversations. Many people who agree with the League position have read postings and taken no action, but have joined later when asked individually.

So the problem is to make it quick and easy to ask people individually. I have developed a method for this.

Most of you are busy studying, doing research or developing software, but you do have time to recruit new members if you follow the efficient method that I have worked out. It requires no special skill at dealing with people—I don’t have any. What it requires is a decision to do the work—even for a few minutes each day.

Here is the method that works for me. When I read a message on a mailing list (or news group) that I can see a way to answer in a helpful tone, I reply to the sender alone (not the entire list), and add on at the end the following text:

By the way, on another topic, what do you think about look and feel lawsuits? Do you think it is a good thing for interfaces to be copyrighted?

I do not mark this as a postscript or as a signature—I feel that would be a suggestion not to give it full attention. (It would be interesting for somebody to verify whether this feeling is accurate.) It is simply the last paragraph of my reply.

Most of the time the person will answer, either with an opinion or with a request for information.

If the person agrees with the League, I send the position papers and membership form, with an introduction such as, “Based on your opinions, it sounds like you agree with the League for Programming Freedom. Perhaps you would like to join; that would be one way in participating in the decision.” I vary this according to what the person said.

If the person wants more information or doesn’t have a position yet, then I send just the look and feel position paper, with a suitable brief introduction explaining what it is about. (For example, “Here is an explanation of what look and feel lawsuits are, and why some of us think they endanger the software field.” This has to vary depending on what the person said.)

If the person claims to see both sides of the issue, then I send a copy of that position paper, introducing it by saying that I hope it will convince him or her that look and feel copyright is purely and simply a bad thing. These people usually state some of the reasons they see, so I respond briefly to the principal points before including the position paper.

Only a few people say they support look and feel lawsuits. I respond briefly to the arguments they give, then refer them to the position paper for a full rebuttal. I introduce the paper with words such as, “Please let me know whether it convinces you that interface copyright is a bad idea.” I don’t spend much time arguing with them; I figure that if they position papers won’t convince them, nothing is likely to. You can’t win ’em all, and we don’t need to.

Occasionally someone asks me if that question is my ‘.signature’. I respond with the truth, which is that I try to ask everybody once. Occasionally I forget and ask somebody twice. Then I apologize.

If I think I might have asked someone already, but I’m not sure, then I phrase the question differently: “Have I asked you yet what you think about look and feel lawsuits?”

It is important to try to avoid taking a tone that says that everyone is obligated to agree with us. That tone makes them less likely to be convinced. But when someone seems to agree but also seems apathetic, sometimes I say, “Please join, if you do agree. This is a big job, so please don’t leave it up to the rest of us. We need your help.”

These methods may not be ideal, but they do seem to work pretty well. So I suggest you start with them and then experiment to find what works best for you.

Using the position papers to do most of the arguing means you don’t have to spend a lot of time rewriting the usual arguments. Also, it means you can supply all the things we have to say on the issue in a single message. The correspondent can then consider it without any feeling of pressure and without your having to do a lot of work on back-and-forth. Most of these discussions involve sending just two or three messages. Most of these people join the League.

I find that this takes me just a few minutes total for one person most of the time. Sometimes it takes ten minutes to write the brief reply when someone does not agree to start with. Occasionally, I waste half an hour trying to rebut when somebody is seriously confused or disagrees, but I think that’s just because I get carried away and forget that there are faster ways to answer. You can probably avoid this problem.

I do this every time I have a chance, sometimes starting ten of these discussions in one day. That takes a lot of time. But if you start just one conversation each day, you will hardly miss the time it takes.

I have a single handy file containing the membership form and position papers (prefaced with a brief explanation of the contents that follow), and I have set up an Emacs abbrev that expands into the “By the way …” question, just to make the process faster.

If you don’t read one message a day from a stranger that you could plausibly reply to, look for people you know that you have not yet discussed the issue with.

The method also works pretty well in person. Make printed copies of the position paper, so that you can simply hand them to a person who doesn’t already agree. Give the person a week or so to read them, and then ask whether he or she agrees with them.

Here is the text at the beginning of my handy information file.

Here is some information about the League for Programming Freedom.

The first two pages are the invitation to join and membership form. We usually print them back-to-back.

The following two pages are the two position papers, in Texinfo format. To format these for printing, you need TEX plus the Texinfo macro package that comes with GNU Emacs. If you just want to read them, you’ll find it easy enough to read them without formatting them.

To get on-line copies of the position papers, send mail to league@prep.ai.mit.edu. You can also ftp them anonymously from prep; look in directory ‘/u/emacs/lpf’. If you cannot send us email, and you joined since the last time we mailed out the position papers, then you can send us snail mail to ask for copies.